A sinusoidal transverse wave traveling in the positive direction of an x-axis has
amplitude of 2.0cm, a wavelength of 10c and a frequency of 400Hz. If the wave equation is of the form
1
Expert's answer
2020-03-17T09:22:02-0400

Assume that

A sinusoidal transverse wave traveling in the positive direction of an x-axis has

amplitude of 2.0cm, a wavelength of 10 cm and a frequency of 400 Hz. If the wave equation is of the form


"\xi=A\sin\omega (t-\frac{x}{v})"


find (1) the velocity of propagation of waves and (2) the maximum velocity of the particles of the medium.


Solution


(1)


"v=\lambda\nu=0.1\cdot 400=40" "m\/s"


(2)


"\frac{d\xi}{dt}=A\omega\cos(t-\frac{x}{v})"


"(\frac{d\xi}{dt})_{max}=A\omega=A\cdot2\pi\nu=0.02\cdot2\cdot3.14\cdot400=50.24" "m\/s"